what is the Price of U.S Coupon Bond issued in May 2018 and the Maturity date is May 2020 , with Par value $30 , yearly Coupon payments $5 ,and Interest Rate 7% ?

1
Expert's answer
2021-12-29T12:30:52-0500

The price is:

"P = 5×\frac{1-(1+0.07)^{-2}} {0.07} + \frac{30} {1.07^2} = 35.24."